# Bridging Wright-Fisher and Moran models

### v1

@@ I Diffusion approximation
Let $x_{t}$ ($t=0,1,...$) be a homogeneous Markov chain in a metric space $E_{N}$ (where $N$ is the population size). Now let $\epsilon_{N}$ be a time scaling such that $\lim\limits_{N\rightarrow+\infty}\epsilon_{N}=0$. Let $x=x_{0}$ be the value of the initial state of the Markov chain. Assuming that the Markov chain has a transition function such that, $\forall x\in E_{N}$,
@@ I.2 Application to simple population genetics models
and we can check that the fourth moment is $\mathcal{O}(1/N^{3})$. This time, the right time scaling is $\epsilon_{N}=1/N^{2}$, and $x_{[t/\epsilon_{N}]}$ converges weakly to the diffusion process with generator
@@ I.3 Model with fixed updated fraction
$\mathbb{E}(\Delta X)$ $=\sum_{k}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\sum_{\Delta X}\Delta X\binom{M}{k+\Delta X}p^{k+\Delta X}(1-p)^{M-\Delta X-k}=\sum_{k}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\left(Mp-k\right)$
@@
$\mathbb{E}(\Delta X^{2})$ $=\sum_{k}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\sum_{\Delta X}\Delta X^{2}\binom{M}{k+\Delta X}p^{k+\Delta X}(1-p)^{M-\Delta X-k}$
@@
$\mathcal{L}_{\rho}=\frac{\rho(2-\rho)}{2}x(1-x)\dfrac{\partial^{2}}{\partial x^{2}}+\alpha\rho x(1-x)\dfrac{\partial}{\partial x}.$ (S17)
II Model with fixed updated fraction: leading non-unit eigenvalue
@@
II.1 Derivation from the Chia and Watterson model and link with the variance of viable offspring
@@
In Ref. [4], Chia and Watterson presented a population genetics model for a population with a fixed size of $N$ individuals. In this model, each individual produces offspring according to a probability law represented by a branching process with generating function $f$. At each discrete step $t$, a number $m$ is drawn from a probability distribution $p_{m}$, and $m$ individuals are sampled from the offspring pool, yielding a total pool of individuals is of size $N+m$. Then, the number $r$ of survivors is drawn from a probability distribution $q_{mr}$, and $r$ individuals are picked from the parents and $N-r$ from the $m$ offspring. The distributions $p_{m}$ and $q_{mr}$ are chosen such that $0\leq r\leq N\text{ and }0\leq N-r\leq m$. For completeness, below we summarize results from [4], and apply them to our partial sampling model.
@@
Application: partial sampling with fixed number of offspring and survivors.
Our partial sampling model can be seen as a particular case of the Chia and Watterson model, where we produce exactly the same number of offspring $M$ at each step, keep those $M$ individuals, and complete the new generation with $N-M$ individuals from the previous generation. This means that $p_{m}=1$ if $m=M$ and 00 otherwise. We also have $q_{mr}=1$ if $r=N-M$, and 0 otherwise. We consider the no-mutation case where $\alpha_{1}=\alpha_{2}=0$. For the branching process representing offspring production, we use the Poisson law $\mathscr{P}(\mu)$ with generating function $f(z)=\exp(\mu(z-1))$. Let us write the transition matrix elements as $p_{ij}$, representing the probability of going from $i$ to $j$ mutants. We have
@@
Chia and Watterson provided an explicit expression of all eigenvalues of the transition matrix for neutral mutants in Ref. [4]. Here, we are interested in the second largest eigenvalue $\lambda_{*}$ (the leading eigenvalue is equal to 1 and has multiplicity 2 when there are no mutations). Applying Chia and Watterson’s formula to our partial sampling model yields
@@
For our partial sampling model, the sums over $m$ and $r$ in the mean lifetime $\mu$ simplify and we get
@@
From Eq. S30, and the expression of $\lambda_{*}$ in Eq. S28, we obtain for our partial sampling model $\phi^{\prime\prime}(1)=2-\rho$. From the diffusion approximation applied to the partial sampling model, and assuming $s\ll 1$ while $Ns\gg 1$ and $N\gg 1$, the fixation probability of one individual can be written as
@@
II.2 Derivation of the leading non-unit eigenvalue using Ref. [2]
Here, we present another method proposed in [2] to compute the leading non-unit eigenvalue $\lambda_{*}$ of the transition matrix for the partial sampling model in the neutral case. The key result is that, if there exists a function $f$ such that $\mathbb{E}[f(X_{t+1})|X_{t}]=af(X_{t})$ where $a$ is a constant, then $a=\lambda_{*}$ is the desired eigenvalue [2]. Taking $f(X)=X(N-X)$ for our partial sampling model yields
@@
$=X_{t}(N-X_{t})\left[1-\frac{\rho(2-\rho)}{N}+o\left(\frac{1}{N}\right)\right],$ (S34)
where we used the fact that $\mathbb{E}(\Delta X)=0$ in the neutral case, and the expression of $\mathbb{E}(\Delta X^{2})$ from Eq. S10. Therefore, we can identify
$\lambda_{*}=1-\frac{\rho(2-\rho)}{N}+o\left(\frac{1}{N}\right),$ (S35)
@@
II.3 Rate of approach to the stationary state
The largest non-unit eigenvalue $\lambda_{*}$ of the transition matrix $\bm{P}=\left\{P_{i\rightarrow j}\right\}_{ij}$ quantifies the rate of convergence towards the stationary state. Concretely, let us consider the vector $\bm{\pi}_{t}=(p_{0}(t),p_{1}(t),p_{2}(t),...,p_{N}(t))$ where $p_{k}(t)$ denotes the probability of having $k$ mutants at time $t$ starting from $1$ mutant at $t=0$. This vector satisfies
@@ II.3 Rate of approach to the stationary state
with $\bm{\pi}_{0}=(0,1,0,0,...,0)$ at time $t=0$. We can decompose this vector in the basis of eigenvectors of $\bm{P}$ written $\bm{u}_{k}$:
@@
where $\bm{\pi}_{\infty}=(1-1/N,0,0,...,0,1/N)$. We can thus write

### v2

@@ I Diffusion approximation
Let $x_{t}$ ($t=0,1,\dots$) be a homogeneous Markov chain in a metric space $E_{N}$ (where $N$ is the population size). Now let $\epsilon_{N}$ be a time scaling such that $\lim\limits_{N\rightarrow+\infty}\epsilon_{N}=0$. Let $x=x_{0}$ be the value of the initial state of the Markov chain. Assuming that the Markov chain has a transition function such that, $\forall x\in E_{N}$,
@@ I.2 Application to simple population genetics models
and we can check that the fourth moment is $\mathcal{O}(1/N^{3})$. This time, we define the time scaling as $\epsilon_{N}=1/N^{2}$, such that $x_{[t/\epsilon_{N}]}$ converges weakly to the diffusion process with generator
@@
Note that we have some freedom in choosing the time scaling (up to a multiplicative constant), and that if another of these scalings was chosen, $x_{[t/\epsilon_{N}]}$ would simply converge to a different diffusion process with a rescaled generator. We choose this definition of $\epsilon_{N}$ to recover the usual diffusion equation for the Moran process. Other choices would just lead to a rescaling of time compared to it.
@@ I.3 Model with fixed updated fraction
$\mathbb{E}(\Delta X)$ $=\sum_{k=0}^{\min(i,M)}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\sum_{\Delta X}\Delta X\binom{M}{k+\Delta X}p^{k+\Delta X}(1-p)^{M-\Delta X-k}=\sum_{k}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\left(Mp-k\right)$
@@
$\mathbb{E}(\Delta X^{2})$ $=\sum_{k=0}^{\min(i,M)}\frac{\binom{i}{k}\binom{N-i}{M-k}}{\binom{N}{M}}\sum_{\Delta X}\Delta X^{2}\binom{M}{k+\Delta X}p^{k+\Delta X}(1-p)^{M-\Delta X-k}$
@@
$\mathcal{L}_{\rho}=\frac{2-\rho}{2}x(1-x)\dfrac{\partial^{2}}{\partial x^{2}}+\alpha x(1-x)\dfrac{\partial}{\partial x}.$ (S17)
Again, note that we have some freedom in choosing the time scaling (up to a multiplicative constant). We define $\epsilon_{N}$ this way to recover the usual Moran and Wright-Fisher diffusion equations in the limiting scenarios. Other choices would just lead to a rescaling of time compared to it.
I.4 Fixation probability in the diffusion approximation
Here, we provide additional plots of the mutant fixation probability in our partial update model, within the diffusion approximation. This fixation probability is given in Eq. 11 in the main text. Figs. S1 and S2 highlight the impact of the fraction $\rho$ of the population that is updated at each discrete time step on the mutant fixation probability.
Figure S1: Fixation probability $\phi_{1}=\phi(1/N)$ (Eq. 11 in the main text) as a function of the updated fraction $\rho$ for different relative fitness advantages $s$ of the mutant. Here $N=100$ and the initial mutant fraction is $x=1/N$.
Figure S2: Fixation probability $\phi(x)$ (Eq. 11 in the main text) as a function of the updated fraction $\rho$ for different initial mutant fractions $x$ with $N=100$.
II Mean absorption times
In this section, we derive the expressions of the mean absorption times. We first consider the unconditional mean absorption time such that either one of the two absorbing barriers $x=0$ and $x=1$ (i.e. mutant extinction or fixation) is reached. Next, we condition the absorption on fixation and on extinction, leading to the mean fixation and extinction times.
II.1 Mean unconditional absorption time
We would like to determine the mean time $t(x)$ it takes to reach mutant extinction or fixation, starting from a given initial mutant fraction $x$. The unconditional mean absorption time $t(x)$, expressed in number of discrete time steps, satisfies the equation [40]
$\mathcal{L}_{\rho}t(x)=-1\,,$ (S18)
with the generator in Eq. S17, and with the boundary conditions $t(0)=t(1)=0$. The integration of this differential equation yields
$t(x)=\frac{1-e^{-2ax/b}}{a(1-e^{-2a/b})}\int_{0}^{1}\mbox{d}y\frac{1-e^{-2a(1-y)/b}}{y(1-y)}-\frac{1}{a}\int_{0}^{x}\mbox{d}y\frac{1-e^{-2a(x-y)/b}}{y(1-y)}\,,$ (S19)
where $a=\rho s$ and $b=\rho(2-\rho)/N$. By splitting the first integral in two, we can write
$t(x)=\frac{1-e^{-2ax/b}}{a(1-e^{-2a/b})}\left\{\int_{0}^{x}\mbox{d}y\frac{1-e^{-2a(1-y)/b}}{y(1-y)}+\int_{x}^{1}\mbox{d}y\frac{1-e^{-2a(1-y)/b}}{y(1-y)}\right\}-\frac{1}{a}\int_{0}^{x}\mbox{d}y\frac{1-e^{-2a(x-y)/b}}{y(1-y)}\,.$ (S20)
Rearranging terms yields
$t(x)=\frac{1-e^{2sN(1-x)/(2-\rho)}}{\rho s\left(1-e^{2sN/(2-\rho)}\right)}\int_{0}^{x}\mbox{d}y\frac{e^{2sNy/(2-\rho)}-1}{y(1-y)}+\frac{1-e^{-2sNx/(2-\rho)}}{\rho s\left(1-e^{-2sN/(2-\rho)}\right)}\int_{x}^{1}\mbox{d}y\frac{1-e^{-2sN(1-y)/(2-\rho)}}{y(1-y)}\,.$ (S21)
This formula is consistent with the general solution $(4.21)$ given in Ref. [40].
II.2 Mean conditional absorption times
We can rewrite Eq. S21 as
$t(x)=\int_{0}^{x}\mbox{d}y~{}t_{1}(x,y)+\int_{x}^{1}\mbox{d}y~{}t_{2}(x,y)\,,$ (S22)
where
$t_{1}(x,y)$ $=\frac{\left(1-e^{2sN(1-x)/(2-\rho)}\right)\left(e^{2sNy/(2-\rho)}-1\right)}{\rho sy(1-y)\left(1-e^{2sN/(2-\rho)}\right)}\,,$ (S23)
$t_{2}(x,y)$ $=\frac{\left(1-e^{-2sNx/(2-\rho)}\right)\left(1-e^{-2sN(1-y)/(2-\rho)}\right)}{\rho sy(1-y)\left(1-e^{-2sN/(2-\rho)}\right)}\,.$ (S24)
Mean mutant fixation time.
The mean absorption time conditioned on mutant fixation can then be expressed as [40]
$t^{*}(x)=\int_{0}^{x}\mbox{d}y~{}t_{1}^{*}(x,y)+\int_{x}^{1}\mbox{d}y~{}t_{2}^{*}(x,y)\,,$ (S25)
with
$t_{i}^{*}(x,y)=t_{i}(x,y)\frac{\phi(y)}{\phi(x)},~{}\textrm{for}\,\,i\in\left\{1,2\right\}.$ (S26)
where $\phi(x)$ is the mutant fixation probability starting from an initial mutant fraction $x$ given in Eq. 11 in the main text. It yields
$t^{*}(x)$ $=\frac{1-e^{2sN(1-x)/(2-\rho)}}{\rho s(1-e^{2sN/(2-\rho)})(1-e^{-2sNx/(2-\rho)})}\int_{0}^{x}\mbox{d}y\frac{\left(1-e^{-2sNy/(2-\rho)}\right)\left(e^{2sNy/(2-\rho)}-1\right)}{y(1-y)}$
$+\frac{1}{\rho s(1-e^{-2sN/(2-\rho)})}\int_{x}^{1}\mbox{d}y\frac{\left(1-e^{-2sN(1-y)/(2-\rho)}\right)\left(1-e^{-2sNy/(2-\rho)}\right)}{y(1-y)}\,,$ (S27)
which satisfies $t^{*}(1)=0$.
In particular, when $s\ll 1$ and $Ns\gg 1$, the leading term of the mean fixation time in Eq. S27 starting from one mutant ($x=1/N$) reads
$t^{*}\left(\frac{1}{N}\right)=\frac{2}{\rho s}\log\left(\frac{2Ns}{2-\rho}\right).$ (S28)
As this time is expressed in number of discrete time steps, in generations we have a mean fixation time
$\rho\,t^{*}\left(\frac{1}{N}\right)=\frac{2}{s}\log\left(\frac{2Ns}{2-\rho}\right)=\frac{2}{s}\log\left(2N_{e}s\right),$ (S29)
where we introduced the effective population size $N_{e}=N/(2-\rho)$, see Section II.6 of the main text. This expression is consistent with the fixation time of a beneficial mutant of order $(1/s)\log(Ns)$ generations found in Ref. [67], where stochastic birth and death events are modeled within a continuous-time branching process.
Mean mutant extinction time.
Similarly, the mean absorption time conditioned on mutant extinction is given by
$t^{**}(x)=\int_{0}^{x}\mbox{d}y~{}t_{1}^{**}(x,y)+\int_{x}^{1}\mbox{d}y~{}t_{2}^{**}(x,y)$ (S30)
with
$t_{i}^{**}(x,y)=t_{i}(x,y)\frac{\psi(y)}{\psi(x)},~{}\textrm{for}\,\,i\in\left\{1,2\right\}.$ (S31)
where $\psi(x)=1-\phi(x)$ is the mutant extinction probability starting from an initial mutant fraction $x$. This gives
$t^{**}(x)$ $=\frac{1}{\rho s(1-e^{2sN/(2-\rho)})}\int_{0}^{x}\mbox{d}y\frac{\left(e^{2sNy/(2-\rho)}-1\right)\left(1-e^{2sN(1-y)/(2-\rho)}\right)}{y(1-y)}$
$+\frac{1-e^{-2sNx/(2-\rho)}}{\rho s(1-e^{-2sN/(2-\rho)})(1-e^{2sN(1-x)/(2-\rho)})}\int_{x}^{1}\mbox{d}y\frac{\left(1-e^{2sN(1-y)/(2-\rho)}\right)\left(1-e^{-2sN(1-y)/(2-\rho)}\right)}{y(1-y)}\,,$ (S32)
which satisfies $t^{**}(0)=0$.
Note that the expressions obtained above satisfy the following symmetry property: $t^{*}(x)=t^{**}(1-x)$. To show this, let us consider
$t^{**}(1-x)$ $=\frac{1}{\rho s(1-e^{2sN/(2-\rho)})}\int_{0}^{1-x}\mbox{d}y\frac{\left(e^{2sNy/(2-\rho)}-1\right)\left(1-e^{2sN(1-y)/(2-\rho)}\right)}{y(1-y)}$
$+\frac{1-e^{-2sN(1-x)/(2-\rho)}}{\rho s(1-e^{-2sN/(2-\rho)})(1-e^{2sNx/(2-\rho)})}\int_{1-x}^{1}\mbox{d}y\frac{\left(1-e^{2sN(1-y)/(2-\rho)}\right)\left(1-e^{-2sN(1-y)/(2-\rho)}\right)}{y(1-y)}$ (S33)
A change of variable gives
$t^{**}(1-x)$ $=\frac{1}{\rho s(1-e^{2sN/(2-\rho)})}\int_{x}^{1}\mbox{d}y\frac{\left(e^{2sN(1-y)/(2-\rho)}-1\right)\left(1-e^{2sNy/(2-\rho)}\right)}{y(1-y)}$
$+\frac{1-e^{-2sN(1-x)/(2-\rho)}}{\rho s(1-e^{-2sN/(2-\rho)})(1-e^{2sNx/(2-\rho)})}\int_{0}^{x}\mbox{d}y\frac{\left(1-e^{2sNy/(2-\rho)}\right)\left(1-e^{-2sNy/(2-\rho)}\right)}{y(1-y)}$ (S34)
$=\frac{1}{\rho s(1-e^{-2sN/(2-\rho)})}\int_{x}^{1}\mbox{d}y\frac{\left(1-e^{-2sN(1-y)/(2-\rho)}\right)\left(1-e^{-2sNy/(2-\rho)}\right)}{y(1-y)}$
$+\frac{1-e^{2sN(1-x)/(2-\rho)}}{\rho s(1-e^{2sN/(2-\rho)})(1-e^{-2sNx/(2-\rho)})}\int_{0}^{x}\mbox{d}y\frac{\left(1-e^{-2sNy/(2-\rho)}\right)\left(e^{2sNy/(2-\rho)}-1\right)}{y(1-y)}=t^{*}(x).$ (S35)
This symmetry is a consequence of the conditioning and of the fact that the operator $\mathcal{L}_{\rho}$ displays the symmetry $x\leftrightarrow 1-x$. Note however that in general $t(x)\neq t(1-x)$, due to the difference of fitness between mutant and wild-type individuals. However, in the neutral case ($s=0$), we find
$t(x)=-\frac{2N}{\rho(2-\rho)}\left[x\log(x)+(1-x)\log(1-x)\right]=-\frac{2N_{e}}{\rho}\left[x\log(x)+(1-x)\log(1-x)\right],$ (S36)
and thus $t(x)=t(1-x)$ for $s=0$.
III Leading non-unit eigenvalue
@@
III.1 Derivation from the Chia and Watterson model and link with the variance of viable offspring
@@
In Ref. [17], Chia and Watterson presented a population genetics model for a population with a fixed size of $N$ individuals. In this model, each individual produces offspring according to a probability law represented by a branching process with generating function $f$. At each discrete step $t$, a number $m$ is drawn from a probability distribution $p_{m}$, and $m$ individuals are sampled from the offspring pool, yielding a total pool of individuals is of size $N+m$. Then, the number $r$ of survivors is drawn from a probability distribution $q_{mr}$, and $r$ individuals are picked from the parents and $N-r$ from the $m$ offspring. The distributions $p_{m}$ and $q_{mr}$ are chosen such that $0\leq r\leq N\text{ and }0\leq N-r\leq m$. For completeness, below we summarize results from [17], and apply them to our partial update model.
@@
Application: partial update with fixed number of offspring and survivors.
Our partial update model can be seen as a particular case of the Chia and Watterson model, where we produce exactly the same number of offspring $M$ at each step, keep those $M$ individuals, and complete the new generation with $N-M$ individuals from the previous generation. This means that $p_{m}=1$ if $m=M$ and 00 otherwise. We also have $q_{mr}=1$ if $r=N-M$, and 0 otherwise. We consider the no-mutation case where $\alpha_{1}=\alpha_{2}=0$. For the branching process representing offspring production, we use the Poisson law $\mathscr{P}(\mu)$ with generating function $f(z)=\exp(\mu(z-1))$. Let us write the transition matrix elements as $p_{ij}$, representing the probability of going from $i$ to $j$ mutants. We have
@@
Chia and Watterson provided an explicit expression of all eigenvalues of the transition matrix for neutral mutants in Ref. [17]. Here, we are interested in the second largest eigenvalue $\lambda_{*}$ (the leading eigenvalue is equal to 1 and has multiplicity 2 when there are no mutations). Applying Chia and Watterson’s formula to our partial update model yields
@@
For our partial update model, the sums over $m$ and $r$ in the mean lifetime $\mu$ simplify and we get
@@
From Eq. S49, and the expression of $\lambda_{*}$ in Eq. S47, we obtain for our partial update model $\phi^{\prime\prime}(1)=2-\rho$. From the diffusion approximation applied to the partial update model, and assuming $s\ll 1$ while $Ns\gg 1$ and $N\gg 1$, the fixation probability of one individual can be written as
@@
III.2 Derivation of the leading non-unit eigenvalue using Ref. [40]
Here, we present another method proposed in [40] to compute the leading non-unit eigenvalue $\lambda_{*}$ of the transition matrix for the partial update model in the neutral case. The key result is that, if there exists a function $f$ such that $\mathbb{E}[f(X_{t+1})|X_{t}]=af(X_{t})$ where $a$ is a constant, then $a=\lambda_{*}$ is the desired eigenvalue [40]. Taking $f(X)=X(N-X)$ for our partial update model yields
@@
$=X_{t}(N-X_{t})\left[1-\frac{\rho(2-\rho)}{N}+o\left(\frac{\rho}{N}\right)\right],$ (S53)
where we used the fact that $\mathbb{E}(\Delta X)=0$ in the neutral case, and the expression of $\mathbb{E}(\Delta X^{2})$ from Eqs. S10 and S15. Therefore, we can identify
$\lambda_{*}=1-\frac{\rho(2-\rho)}{N}+o\left(\frac{\rho}{N}\right),$ (S54)
@@
III.3 Rate of approach to the stationary state
The largest non-unit eigenvalue $\lambda_{*}$ of the transition matrix $\bm{P}=\left\{P_{i\rightarrow j}\right\}_{ij}$ quantifies the rate of convergence towards the stationary state. Concretely, let us consider the vector $\bm{\pi}_{t}=(p_{0}(t),p_{1}(t),p_{2}(t),\dots,p_{N}(t))$ where $p_{k}(t)$ denotes the probability of having $k$ mutants at time $t$ starting from $1$ mutant at $t=0$. This vector satisfies
@@ III.3 Rate of approach to the stationary state
with $\bm{\pi}_{0}=(0,1,0,0,\dots,0)$ at time $t=0$. We can decompose this vector in the basis of eigenvectors of $\bm{P}$ written $\bm{u}_{k}$:
@@
where $\bm{\pi}_{\infty}=(1-1/N,0,0,\dots,0,1/N)$. We can thus write